# MOON: Multi-Objective Optimization-Driven Object-Goal Navigation Using a Variable-Horizon Set-Orienteering Planner

### v1

@@ I Introduction
Despite these advances, we observe that efforts to leverage the long-standing knowledge of the optimization techniques community—especially optimal planning techniques [7]—in ON methods are either overlooked or remain underdeveloped. Consequently, many systems fall into myopic behaviors and fail to balance exploratory and goal-directed actions effectively. While existing ON methods often encourage visiting semantically relevant landmarks, they largely neglect strategic planning of the visiting order and global path optimization, thereby underutilizing valuable insights from the optimization community. This issue becomes critical in real-world environments, which are often large-scale and complex, yet many existing methods rely on idealized simulators such as Habitat-Sim [8] and AI2-THOR [9], which oversimplify sensor models and spatial complexity, limiting real-world applicability. Furthermore, in multi-robot scenarios, landmark maps can be shared to build a richer global knowledge base, but effective utilization of such shared information for coordinated strategic navigation remains underexplored. We argue that integrating intuitive, myopic high-level reasoning from existing TFP/RLP/ZSP frameworks with globally optimal planning is essential for effective and efficient ON in large, complex real-world settings.
@@
In this work, we reformulate the standard single-target ON task as a multi-objective optimization-driven object goal navigation (MOON), and as a case study, model two conflicting objectives: frontier exploration for discovering unknown landmarks and efficient touring of previously observed landmarks. Our formulation emphasizes the intrinsic multi-objective nature of ON tasks and is both orthogonal and complementary to existing problem settings, such as multi-ON tasks [10], which address multiple target objects in an optionally specified order, and general multi-objective navigation [11]. Specifically, we model ON planning as a variable-horizon optimization problem aiming to maximize average SPL (Success weighted by Path Length [12]) based on incrementally constructed landmark maps. Although our method is compatible with the TFP, RLP, and ZSP frameworks, we adopt the TFP framework from [1] as the basis for our experiments. In this framework, the robot detects objects from RGBD and semantic images and semantically estimates landmark objects related to the target. We then extend TFP by introducing a query-based occupancy map (QOM) [13], which compactly describes essential landmark information by discarding data irrelevant to object-target navigation and focusing solely on relevant landmarks. Finally, we formulate the landmark visiting order optimization problem (LOOP) as a set orienteering problem (SOP) and further extend it to a variable-horizon SOP that incorporates map updates. Notably, the proposed method facilitates map sharing and multi-objective navigation, enhancing the efficiency and scalability of ON tasks by integrating semantic reasoning with optimal planning.
@@ I-A Relation to Prior Work
The QOM-based object-goal navigation (ON) approach employed in this research is not new and was developed around 2005 by researchers at the Mathematical Engineering Laboratory, Department of Mechanical Engineering, Graduate School of Kyushu University [13]. At that time, an active vision task that was conceptually equivalent to ON was studied. However, unlike the now-common ON setup, which involves exploration of the physical real-world environment, the robot’s workspace in that earlier research was its internal memory space—namely, visual experiences—rather than the physical environment itself, a context in which offline reinforcement learning [15] plays a critical role. Several follow-up studies have since been conducted.
@@
In recent years, we have independently developed a series of methods aligned with the modern ON setup, including TFP [17], RLP [18], and ZSP [19].
@@ II System Overview
This section describes our method that replaces the myopic TFP/RLP/ZSP in conventional ON systems with a multi-objective optimization planner that guarantees global optimality. We begin by introducing our extended problem setting, which leverages a variable-horizon incremental landmark map (II-A), to clarify the motivation behind our approach. We then present the system design to address this problem. Specifically, we first review the TFP planner proposed in [1] (II-B), then explain how it is extended to a QOM-based planner following the ALCON framework [18] (II-C), and finally, we introduce our proposed multi-objective optimization planner (II-D).
@@ II-A Motivation: Variable-Horizon Incremental Map (VHIM)
With recent advances in ON and AI technologies, there is a growing demand to extend ON systems using variable-horizon incremental landmark maps that allow the mapping horizon to dynamically change during incremental mapping, in contrast to the fixed or small-scale maps typically used in conventional ON tasks. In particular, the following scenarios highlight the emerging significance of such maps:
@@
High-elevation view images that overlook the entire workspace are becoming practical as landmark maps for ON [14].
@@
In multi-objective navigation scenarios, pre-built maps can serve directly as a prior landmark map during ON tasks [11].
@@
In multi-robot coordination scenarios, aerial images captured by a partner drone serve as an alternative landmark map for ON [20].
@@
A wide-area image taken from a distant viewpoint before the robot enters the workspace is also utilized as a landmark map [21].
@@ II-C QOM-based ALCON Planner
The Query-Based Occupancy Map (QOM) [13] extends the conventional concept of occupancy maps [22] from representing the existence probability of objects in each grid cell to representing the existence probability of a specific target object. In the original formulation, QOM leveraged the advanced visual recognition capabilities of a remote human operator to estimate the likelihood of target presence, by training and repeatedly employing a machine learning model (e.g., SVM) that imitates human operators. In contrast, the present study replaces human vision with modern computer vision techniques, enabling autonomous inference.
@@
Recently, QOM was exployed in [17] as a compact yet informative grid map that records the probability of the target object’s presence, and in the ALCON framework in [18], QOM was further utilized to construct a scalable navigation route graph. Specifically, the spatial scene graph used in StructNav is replaced in ALCON with a globally optimal route graph. Path planning is then performed on this graph using an actor-critic reinforcement learning (RL) planner. In this context, a natural implementation of the QOM involves evaluating how relevant each cluster in the semantic point cloud is to the target object. Following StructNav [1], this relevance can be estimated using BERT embeddings and cosine similarity. The proposed ALCON framework [18], while building on the foundation of StructNav, is notable for replacing the myopic frontier-based planner in StructNav with an RL-based planner.
@@ II-D Extension in the MOON framework
In the current study, we introduce a solver for variable-horizon set-orienteering problem (VH-SOP) equipped with a competence module that supervises the planning capabilities of previously developed systems—TFP (Training-Free Planner) [17], RLP (Reinforcement Learning-based Planner) [18], and ZSP (Zero-Shot Planner) [19]—with the goal of achieving globally-optimal path-planning.
@@
To formulate the Set Orienteering Problem (SOP), we explicitly distinguish between two types of visual detection models: a long-range landmark detector with a sensing range (e.g., m), and a short-range target detector with a sensing range (e.g., m). The long-range detector can identify semantically relevant landmarks (e.g., kitchen, desk) within its wide sensing range , and as increases, the number of candidate landmarks also grows, thereby escalating the complexity of planning. In contrast, the short-range detector is capable of recognizing the target object (e.g., a specific mug) only within its limited range , which serves as a crucial constraint in determining whether a given viewpoint qualifies as a valid landmark visit.
@@ III-B Extension to General-Purpose MOON
III-C Evaluation
III-C1 Problem Setting
We consider a large-scale exploration problem in a workspace of size , which is significantly larger than typical simulation environments such as Habitat-Sim. The workspace consists of multiple rooms, each having entrances and exits on four sides.
A single target is placed in the environment. While this work assumes a single target, it can be readily extended to a multi-target setting (multi-ON problem).
@@ III-C1 Problem Setting
III-D Hyperparameters
User-defined Parameters
@@ III-D Hyperparameters
Randomized for Each Trial
@@
III-D1 Implementation Details
Navigation Map
Before navigation begins, the map is initialized as a blank map. During navigation, landmarks and obstacles (e.g., walls) detected within the long-range sensor radius are incrementally added to the map. Redundant entries are ignored, meaning that we assume zero information gain from repeated observations of the same landmark.
Graph Map Construction
Nodes: Each node is defined as a region of radius centered at an observed landmark. Previously observed landmarks are ignored.
@@ III-D1 Implementation Details
III-D2 Evaluation Metric
@@ III-D3 Baselines
For the second planner, we experimented with multiple different implementations. In conclusion, both the second and third planners were found to be unsuitable for direct use in the current task, as they aim to sweep the entire floor area with the robot body and thus tend to generate overly dense paths. Of course, trying different spatial discretizations to generate graph nodes may still be effective and remains a topic for future work. The first and fourth planners, on the other hand, demonstrated meaningful SPL performance. The results are as follows:

### v2

@@ I Introduction
Despite these advances, we observe that efforts to leverage the long-standing knowledge of the optimization techniques community—especially optimal planning techniques [7]—in ON methods are either overlooked or remain underdeveloped (Fig. 1). Consequently, many systems fall into myopic behaviors and fail to balance exploratory and goal-directed actions effectively. While existing ON methods often encourage visiting semantically relevant landmarks, they largely neglect strategic planning of the visiting order and global path optimization, thereby underutilizing valuable insights from the optimization community. This issue becomes critical in real-world environments, which are often large-scale and complex, yet many existing methods rely on idealized simulators such as Habitat-Sim [8] and AI2-THOR [9], which oversimplify sensor models and spatial complexity, limiting real-world applicability. Furthermore, in multi-robot scenarios, landmark maps can be shared to build a richer global knowledge base, but effective utilization of such shared information for coordinated strategic navigation remains underexplored. We argue that integrating intuitive, myopic high-level reasoning from existing TFP/RLP/ZSP frameworks with globally optimal planning is essential for effective and efficient ON in large, complex real-world settings.
@@
In this work, we reformulate the standard single-target ON task as a multi-objective optimization-driven object goal navigation (MOON), and as a case study, model two conflicting objectives: frontier exploration for discovering unknown landmarks and efficient touring of previously observed landmarks (Fig. 2)111 The ”moon and star” representation was introduced as a schematic or cartoon-like abstraction to emphasize visual clarity while maintaining mathematical tractability. Nevertheless, our framework is not limited to point landmarks or circular sensing ranges and can be generalized to a wide range of application domains. . Our formulation emphasizes the intrinsic multi-objective nature of ON tasks and is both orthogonal and complementary to existing problem settings, such as multi-ON tasks [10], which address multiple target objects in an optionally specified order, and general multi-objective navigation [11]. Specifically, we model ON planning as a variable-horizon optimization problem aiming to maximize average SPL (Success weighted by Path Length [12]) based on incrementally constructed landmark maps. Although our method is compatible with the TFP, RLP, and ZSP frameworks, we adopt the TFP framework from [1] as the basis for our experiments. In this framework, the robot detects objects from RGBD and semantic images and semantically estimates landmark objects related to the target. We then extend TFP by introducing a query-based occupancy map (QOM) [13], which compactly describes essential landmark information by discarding data irrelevant to object-target navigation and focusing solely on relevant landmarks. Finally, we formulate the landmark visiting order optimization problem (LOOP) as a set orienteering problem (SOP) and further extend it to a variable-horizon SOP that incorporates map updates. Notably, the proposed method facilitates map sharing and multi-objective navigation, enhancing the efficiency and scalability of ON tasks by integrating semantic reasoning with optimal planning.
@@ I-A Relation to Prior Work
The QOM-based object-goal navigation (ON) approach employed in the current research is not new and was developed around 2005 by researchers at the Mathematical Engineering Laboratory, Department of Mechanical Engineering, Graduate School of Kyushu University [13] (QOM [IROS05]). At that time, an active vision task that was conceptually equivalent to ON was studied. However, unlike the now-common ON setup, which involves exploration of the physical real-world environment, the robot’s workspace in that earlier research was its internal memory space—namely, visual experiences—rather than the physical environment itself, a context in which offline reinforcement learning [15] plays a critical role. Several follow-up studies have since been conducted.
@@
In recent years, we have independently developed a series of methods aligned with the modern ON setup, including TFP [17] (CON [CoRR24a]), RLP [18] (ALCON [CoRR24b]), and ZSP [19] (LGR [CoRR25]).
@@ II System Overview
This section describes our method that oversees the myopic TFP/RLP/ZSP in conventional ON systems by employing a multi-objective optimization planner that guarantees global optimality (Fig. 3). We begin by introducing our extended problem setting, which leverages a variable-horizon incremental landmark map (VHIM), to clarify the motivation behind our approach (II-A). We then present the system design to address this problem. Specifically, we first review the TFP planner proposed in [1] (II-B), then explain how it is extended to a QOM-based planner following the ALCON framework [18] (II-C), and finally, we introduce our proposed multi-objective optimization planner (II-D).
@@ II-A Motivation: Variable-Horizon Incremental Map (VHIM)
With recent advances in ON and AI technologies, there is a growing demand to extend ON systems using variable-horizon incremental landmark maps (VHIMs) that allow the mapping horizon to dynamically change during incremental mapping (Fig. 4), in contrast to the fixed or small-scale maps typically used in conventional ON tasks. In particular, the following scenarios highlight the emerging significance of such maps:
@@ II-C QOM-based ALCON Planner
The Query-Based Occupancy Map (QOM) [13] extends the conventional concept of occupancy maps [24] from representing the existence probability of objects in each grid cell to representing the existence probability of a specific target object (Fig. 6). In the original formulation, QOM leveraged the advanced visual recognition capabilities of a remote human operator to estimate the likelihood of target presence, by training and repeatedly employing a machine learning model (e.g., SVM) that imitates human operators. In contrast, the present study replaces human vision with modern computer vision techniques, enabling autonomous inference.
@@
Recently, QOM was employed in [17] as a compact yet informative grid map that records the probability of the target object’s presence, and in the ALCON framework in [18], QOM was further utilized to construct a scalable navigation route graph. Specifically, the spatial scene graph used in StructNav is replaced in ALCON with a globally optimal route graph. Path planning is then performed on this graph using an actor-critic reinforcement learning (RL) planner. In this context, a natural implementation of the QOM involves evaluating how relevant each cluster in the semantic point cloud is to the target object. Following StructNav [1], this relevance can be estimated using BERT embeddings and cosine similarity. The proposed ALCON framework [18], while building on the foundation of StructNav, is notable for replacing the myopic frontier-based planner in StructNav with an RL-based planner.
@@ II-D Extension in the MOON framework
In the current study, we introduce a solver for the variable-horizon set-orienteering problem (VH-SOP) that has the competence to supervise previously developed systems—TFP (Training-Free Planner) [17], RLP (Reinforcement Learning-based Planner) [18], and ZSP (Zero-Shot Planner) [19]—with the goal of achieving globally-optimal path-planning.
@@
To formulate the Set Orienteering Problem (SOP), we explicitly distinguish between two types of visual recognition models: a long-range landmark detector with a sensing range (e.g., m), and a short-range target detector with a sensing range (e.g., m). The long-range detector can identify semantically relevant landmarks (e.g., kitchen, desk) within its wide sensing range , and as increases, the number of candidate landmarks also grows, thereby escalating the complexity of planning. In contrast, the short-range detector is capable of recognizing the target object (e.g., a specific mug) only within its limited range , which serves as a crucial constraint in determining whether a given viewpoint qualifies as a valid landmark visit.
@@ III-B Extension to General-Purpose MOON
IV Evaluation
IV-A Problem Setting
We consider a large-scale exploration problem in a workspace of size , which is significantly larger than typical simulation environments such as Habitat-Sim. The workspace consists of multiple rooms, each having entrances and exits on four sides (see Fig. 7).
A single target point object is placed in the environment. While this work assumes a single target, it can be readily extended to a multi-target setting (multi-ON problem).
@@ IV-A Problem Setting
IV-B Hyperparameters
IV-B1 User-defined Parameters
@@
Number of trials 100
IV-B2 Randomized for Each Trial
@@
IV-C Implementation Details
IV-C1 Object Grid Map
Before navigation begins, an object grid map is initialized as a blank map. During navigation, landmarks and obstacles (e.g., walls) detected within the long-range sensor radius are incrementally added to the map. Redundant entries are ignored, meaning that we assume zero information gain from repeated observations of the same landmark.
IV-C2 Graph Map Construction
Nodes: Each node is defined as a region of radius centered at an observed landmark. Previously observed landmarks are excluded from the navigation graph.
@@ IV-C2 Graph Map Construction
IV-C3 MOO Solver
The SOP solver from [33] is employed as the MOO solver of our system, and we have confirmed that it works efficiently with the navigation graph recognition module and the action planner module. This paper proposes a new Integer Linear Programming (ILP) formulation for small to medium-sized problems, alongside a Variable Neighborhood Search (VNS) metaheuristic-based algorithm (VNS-SOP) that helps reduce computation time. Beyond offering some improvements to existing benchmark results, it also demonstrates the applicability of its approach to other Orienteering Problem (OP) variants, such as the Dubins Orienteering Problem (DOP) and the Orienteering Problem with Neighborhoods (OPN), suggesting valuable insights.
IV-D Evaluation Metric
@@ IV-E Baselines
For the second planner, we experimented with multiple different implementations.
IV-F Results
In conclusion, both the second and third planners were found to be unsuitable for direct use in the current task, as they aim to sweep the entire floor area with the robot body and thus tend to generate overly dense paths. Of course, trying different spatial discretizations to generate graph nodes may still be effective and remains a topic for future work. The first and fourth planners, on the other hand, demonstrated meaningful SPL performance. The results are as follows: